# Analysis and Synthesis Denoisers

### v2

@@ 5 Numerical experiments
In Section 5.1 we study the stability of the FB-PnP iterations described in Section 2 and Section 3. We in particular aim to illustrate results presented in Theorem 2.5 and in Theorem 3.8, for the analysis and synthesis denoisers (AD and SD, respectively). This study is conducted on a toy compressed sensing example.
@@ 5.1 Stability analysis of the FB-PnP iterations
In this section we aim to analyze the stability of the FB-PnP iterations when using the proposed AD or SD, depending on the number of layers $L$. Specifically, we want to illustrate the results presented in Theorem 2.5 and in Theorem 3.8. Although these results only hold when $L=1$ of $L\to\infty$ (i.e., the proximity operators are computed accurately), we will also empirically investigate the behavior of the AD and SD when $1<L<+\infty$. To this aim, we consider a toy example consisting of an equation system of the form of $y=A\overline{x}$, where $\overline{x}\in\mathbb{R}^{N}$ is a vector of dimension $N=50$ generated randomly following a uniform distribution, $A\colon\mathbb{R}^{N}\to\mathbb{R}^{M}$ is a Gaussian measurement operator with $M=20$, where coefficients are generated randomly following a Gaussian distribution with mean 00 and standard deviation $1$. Both the analysis and synthesis dictionaries are generated as Gaussian dictionaries with mean 00 and standard deviation $1$, of size $\Gamma\in\mathbb{R}^{100\times 50}$ and $D\in\mathbb{R}^{50\times 100}$, respectively.
@@ 5.1.1 Analysis denoiser
Figure 1: Convergence profile of algorithm (2.7) with $L\in\{1,20,50,100\}$. Solutions $(x^{*},u^{*})$ are pre-computed considering algorithm (2.7) with $L\to\infty$. The dictionary $\Gamma$ is learned with $L=1$.
@@
In fig. 1 we show the behavior of $(\|x_{k}-x^{*}\|)_{k\in\mathbb{N}}$ and $(\|u_{k}-u^{*}\|)_{k\in\mathbb{N}}$ with $(x_{k},u_{k})_{k\in\mathbb{N}}$ generated by algorithm (2.7) with $L\in\{1,20,50,100\}$, and $(x^{*},u^{*})$ generated by algorithm (2.7) with $L\to\infty$ (obtained after $10^{4}$ iterations). Note that $x^{*}$ is solution to problem (2.6). We see that the trajectory per iteration is the same independently of the value of $L$. Hence, according to Theorem 2.5 we can deduce that $(x^{*},\tau^{-1}u^{*})$ is solution to the saddle-point problem (2.9). We can further observe that the convergence is faster in time for smaller values of $L$.
@@ 5.1.2 Synthesis denoiser
We shown in Theorem 3.8 that taking $L=1$ in algorithm (3.20) enables solving (3.6) as well as the direct sparse coding problem (3.7). Since algorithm (3.20) for $L\to\infty$ solves (3.6) (theorem 3.7), we deduce that both algorithm (3.20) for $L\to\infty$ and algorithm (3.21) (i.e., $L=1$) solve the same problem (3.6). In this section we aim to illustrate this result by comparing the output of algorithm (3.21) with the output of algorithm (3.20) for $L=10^{4}$. Furthermore, although the results from Proposition 3.1 only hold for $L=1$, we show empirically that they remain true for $1<L<+\infty$ for our problem by considering $L\in\{20,50,100\}$.
In fig. 2 we show the behavior of $(\|x_{k}-x^{*}\|)_{k\in\mathbb{N}}$ and $(\|z_{k}-z^{*}\|)_{k\in\mathbb{N}}$ with $(x_{k},z_{k})_{k\in\mathbb{N}}$ generated by algorithm (3.20) with $L\in\{1,20,50,100\}$, and $(x^{*},z^{*})$ generated by algorithm (3.20) with $L\to\infty$ (obtained after $10^{4}$ iterations). Note that $x^{*}$ is solution to problem (3.6). We see that the trajectory per iteration is fairly the same when $L\in\{20,50,1000\}$, while $L=1$ is taking a bit more iterations to reach convergence. However, for the convergence speed, all strategies seems to behave similarly, with $L\in\{1,20\}$ being very slightly faster.
Figure 2: Convergence profile of algorithm (3.20) with $L\in\{1,20,50,100\}$. Solutions $(x^{*},u^{*})$ are pre-computed considering algorithm (3.20) with $L\to\infty$. The dictionary $D$ is learned with $L=1$.
@@ 5.2 Application to image deblurring within a DDL framework
Although the results presented in this work are generic and can be applied for any linear operators $\Gamma$ and $D$, they are of particular interest in the DDL framework, where these operators are learned. We will explore this case in this section, and apply the resulting methods to an image deblurring problem. For the sake of completeness we will compare the resulting algorithms to a FB-PnP algorithm where the denoiser is a DRUnet.
@@ 5.2.1 DDL context and training
$v=\overline{x}+\upsilon w,$ (5.1)
where $w\in\mathbb{R}^{N}$ is a realization of an i.i.d. standard normal random variable, and $\upsilon>0$ is the scale of the noise. Both analysis and synthesis formulations to find the estimate of $\overline{x}$ from $v$ in (5.1) (see (2.5) and (3.1), respectively) have been extensively studied in the literature (see e.g. [42, 55] for the analysis, and [40, 2, 36] for the synthesis formulation). In the context of DDL, these problems can be reformulated as a joint estimation of the image and the dictionary, or more generally as bi-level optimization problems [36]. Then, the dictionaries are defined as solution to
@@
If in (5.2) (resp. (5.3)), the loss function $\ell^{\mathcal{A}}(\Gamma)$ is the same as the one minimized in (2.5) (resp. if $\ell^{\mathcal{S}}(D)$ is the same as the one minimized in (3.1)), then the bi-level problem is equivalent to the DDL joint estimation problem. This choice is often used in an unsupervised setting. In a supervised setting, where a ground truth dataset can be used, a standard choice for $\ell^{\mathcal{A}}$ and $\ell^{\mathcal{S}}$ is the mean square error (MSE) loss, minimizing the $\ell^{2}$ norm of the difference between the ground truth and the estimated signal $x^{\dagger}_{\Gamma}$ or $x^{\dagger}_{D}$. In the remainder, we will use this approach. Specifically, we aim to learn $\Gamma$ and $D$ such that
$\begin{cases}x^{\dagger}_{\mathcal{A}}=G^{\mathcal{A}}_{L,\upsilon,v}(\Gamma^{*}v)\approx\overline{x},\\ x^{\dagger}_{\mathcal{S}}=G^{\mathcal{S}}_{L,\upsilon,v}(D^{*}v)\approx\overline{x}\end{cases}$ (5.4)
@@
where $G^{\mathcal{A}}$ and $G^{\mathcal{S}}$ are defined in Model 2.2 and Model 3.5, respectively, for $L\geq 1$. Furthermore, we assume that the dictionaries $\Gamma$ and $D$ are parametrized by learnable parameters $\theta_{\Gamma}$ and $\theta_{D}$, respectively (i.e., convolution kernel). Then, given a dataset of pairs of clean/noisy images ${(\overline{x}_{i},v_{i})}_{i\in\mathbb{S}_{\mathcal{T}}}$ obtained following (5.1), the unrolled denoisers in (5.4) are trained by solving, respectively
@@
$\underset{\theta_{\Gamma}}{\operatorname{minimize}}\;\frac{1}{\#(\mathbb{S}_{\mathcal{T}})}\sum_{i\in\mathbb{S}_{\mathcal{T}}}\|G^{\mathcal{A}}_{L_{\text{train}},\upsilon_{i},v_{i}}(\Gamma^{*}v_{i})-\overline{x}_{i}\|,$ (5.5)
and
$\underset{\theta_{D}}{\operatorname{minimize}}\;\frac{1}{\#(\mathbb{S}_{\mathcal{T}})}\sum_{i\in\mathbb{S}_{\mathcal{T}}}\|G^{\mathcal{S}}_{L_{\text{train}},\upsilon_{i},v_{i}}(0_{S})-\overline{x}_{i}\|,$ (5.6)
where $0_{S}\in\mathbb{R}^{S}$ is a constant vector with 00 value on all coefficients. We trained our models on patches of size $150\times 150$ of images from the training dataset from the BSDS500 image bank and we chose a noise level for the training of $\upsilon=0.05$. The learnable parameters $\theta_{\Gamma}$ and $\theta_{D}$ correspond to $50$ convolutional filters of dimension $5\times 5$, and we fix $\sigma=\frac{1.8}{||\Gamma||^{2}}$ in Model 2.2 and $\zeta=\frac{1.8}{||D||^{2}}$ in Model 3.5. We tested two strategies: either training the denoisers with only $L_{\text{train}}=1$ iteration or with $L_{\text{train}}=20$ iterations that share the same dictionary.
@@
In Table 1 we give the inference time for denoising an image, when applying the AD and SD with either $L_{\text{den}}=1$ or $L_{\text{den}}=20$ iterations. For comparison, we also provide the inference time when using a DRUnet, where the weights have been taken from the DeepInv PyTorch library. We further provide examples of output images for the different denoisers in Figure 3, evaluated in different configurations. It can be observed that the cases where SD is evaluated with $L_{\text{den}}=1$ iteration (SD train $L_{\text{train}}=1,20$ and $L_{\text{den}}=1$) lead to very poor results. When SD is trained with $L_{\text{train}}=1$ and evaluated with $L_{\text{den}}=20$ iterations, results start to improve, although remaining quite poor. Additional simulations have shown that increasing $L_{\text{den}}$ in the evaluation continues improving the output. All other cases look very similar, with DRUnet giving best denoising results.
@@
Table 1: Average time and standard deviation when applying different denoisers, computed over $150$ images from BSDS500 cropped to $150\times 150$, with noise level $\upsilon=0.05$.
@@
Averaged time ($\pm$±std) $0.19(\pm 0.08)$ $3.94(\pm 1.46)$ $0.27(\pm 0.16)$ $3.12(\pm 0.96)$ $5.32(\pm 0.76)$
(×$\times 10^{-2}$ sec.)
@@
$\overline{x}$ $v$ $L_{\text{den}}=1$ $L_{\text{den}}=20$ $L_{\text{den}}=1$ $L_{\text{den}}=20$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
DRUnet $L_{\text{den}}=1$ $L_{\text{den}}=20$ $L_{\text{den}}=1$ $L_{\text{den}}=20$
@@
Figure 3: Denoising result example for different denoisers and configurations. Top left two images show original image $\overline{x}$ and noisy image $v$. Bottom left image shows denoised image with DRUnet. Central four images show SD (top) and AD (bottom) trained with only $L_{\text{train}}=1$ iteration, and evaluated with either $L_{\text{den}}=1$ (left) or $L_{\text{den}}=20$ (rigth) iterations for denoising. Right four images show SD (top) and AD (bottom) trained with $L_{\text{train}}=20$ iteration, and evaluated with either $L_{\text{den}}=1$ (left) or $L_{\text{den}}=20$ (rigth) iterations for denoising.
@@ 5.2.2 Simulation results
blur kernel
     ground truth observation ground truth observation Refer to caption Refer to caption   Refer to caption Refer to caption
@@
We consider an image deblurring problem of the form of (1.1), where $A\colon\mathbb{R}^{N}\to\mathbb{R}^{M}$ models a convolution with kernel given in fig. 4 and $\varepsilon=0.05$. Both the AD and SD are trained following the DDL procedure described in section 5.2.1. The resulting AD and SD are then plugged in the FB-PnP iterations and applied to the image deblurring problem. We will further compare the proposed approaches with an FB-PnP algorithm with a DRUnet denoiser [56], considered as state-of-the-art for PnP.
                                 $\lambda=10^{-5}$                                      $\lambda=10^{-3}$                                        $\lambda=10^{-2}$                                      $\lambda=10^{-1}$
SD train $L_{\text{train}}=1$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
SD train $L_{\text{train}}=20$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
AD train $L=1$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
AD train $L=20$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
DRUnet
@@
Figure 5: Convergence profiles with respect to iterations for the different FB-PnP algorithms. AD and SD are trained with either $L_{\text{train}}=1$ or $L_{\text{train}}=20$, and implemented in the FB-PnP with either $L_{\text{PnP}}=1$ or $L_{\text{PnP}}=20$.
                                 $\lambda=10^{-5}$                                      $\lambda=10^{-3}$                                        $\lambda=10^{-2}$                                      $\lambda=10^{-1}$
SD train $L=1$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
SD train $L=20$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
AD train $L=1$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
AD train $L=20$
$L_{\text{PnP}}=1$
$L_{\text{PnP}}=20$
DRUnet
@@
Figure 6: PSNR profiles with respect to iterations for the different FB-PnP algorithms. AD and SD are trained with either $L_{\text{train}}=1$ or $L_{\text{train}}=20$, and implemented in the FB-PnP with either $L_{\text{PnP}}=1$ or $L_{\text{PnP}}=20$
@@
In Figure 5, we compare convergence behaviors of the FB-PnP iterations $(\|x_{k+1}-x_{k}\|)_{k\in\mathbb{N}}$ for different denoisers, for different regularization parameters $\lambda\in\{10^{-5},10^{-3},10^{-2},10^{-1}\}$. Each plot shows convergence curves when running the algorithm on $5$ different images. The first two rows correspond to SD trained with $L_{\text{train}}=1$, second two rows show SD trained with $L_{\text{train}}=20$, third two rows show AD trained with $L_{\text{train}}=1$ and fourth two rows show AD trained with $L_{\text{train}}=20$. In each case we run the FB-PnP with either $L_{\text{PnP}}=1$ or $L_{\text{PnP}}=20$ sub-iterations. The very last row in Figure 5 shows FB-PnP behaviour with DRUnet. FB-PnP algorithms with SD and AD exhibit monotonic convergence trends independently of the choice of the regularization parameter $\lambda$. This is not the case for DRUnet that seems to converge only for $\lambda=10^{-1}$, while other choices are not stable [^1] For DRUnet $\lambda$ corresponds to the noise level given as an input to the network. Additional simulations have been performed for DRUnet increasing $\lambda$, however these choices were observed to lead to lower PSNR values.. We further provide curves of PSNR values with respect to iterations in Figure 6, for the above mentioned FB-PnP iterations. It can be observed that for AD the PSNR profiles are very similar when fixing $\lambda$, and looking at the different strategies $(L_{\text{train}},L_{\text{PnP}})\in\{1,20\}^{2}$. This suggests that the AD framework is very stable, independently on the number of sub-iterations used for training or for the reconstruction process. This is not the case for SD. We can observe that the strategies $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ and $(L_{\text{train}},L_{\text{PnP}})=(20,20)$ have similar trends, while strategies $(L_{\text{train}},L_{\text{PnP}})=(1,1)$ and $(L_{\text{train}},L_{\text{PnP}})=(20,1)$ are much slower. This was to be expected given the denoising performances of the last two strategies, shown in Figure 3 for $(L_{\text{train}},L_{\text{den}})=(1,1)$ and $(L_{\text{train}},L_{\text{den}})=(20,1)$. DRUnet exhibit very unstable behaviour for $\lambda\in\{10-5,10^{-3}\}$, as expected given Figure 5. More stable results are for $\lambda=10^{-1}$.
SD with $L_{\text{train}}=1$ SD with $L_{\text{train}}=20$
$\overline{x}$ $y$ $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$ $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PSNR=$27.43$dB PSNR=$25.71$dB PSNR=$27.38$dB PSNR=$22.13$dB
AD with $L_{\text{train}}=1$ AD with $L_{\text{train}}=20$
DRUnet $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$ $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$
PSNR=$27.07$dB PSNR=$29.06$dB PSNR=$29.06$dB PSNR=$28.09$dB PSNR=$28.09$dB
SD with $L_{\text{train}}=1$ SD with $L_{\text{train}}=20$
$\overline{x}$ $y$ $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$ $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PSNR=$26.80$dB PSNR=$23.30$dB PSNR=$26.21$dB PSNR=$20.82$dB
AD with $L_{\text{train}}=1$ AD with $L_{\text{train}}=20$
DRUnet $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$ $L_{\text{PnP}}=1$ $L_{\text{PnP}}=20$
PSNR=$25.13$dB PSNR=$27.42$dB PSNR=$27.42$dB PSNR=$26.45$dB PSNR=$26.45$dB
@@
Figure 7: Examples of FB-PnP outputs obtained with the different denoisers tested, for two image examples. Regularization parameters are chosen according to convergence profiles given in Figure 5 and Figure 6: $\lambda_{\textsf{SD}}=10^{-2}$, $\lambda_{\textsf{AD}}=10^{-3}$ and $\lambda_{\textsf{DRUnet}}=10^{-1}$.

### v3

@@ 5 Numerical experiments
In Section 5.1 we study the stability of the FB-PnP iterations described in Section 2 and Section 3. In particular, we aim to illustrate results presented in Theorem 2.5 and in Theorem 3.8, for the analysis and synthesis denoisers (AD and SD, respectively). This study is conducted on a toy compressed sensing example.
@@
Codes to reproduce the results presented in this section are available on GitHub.
@@ 5.1 Stability analysis of the FB-PnP iterations
In this section we consider a simple compressive sensing example, and we aim to analyze the stability of the FB-PnP iterations when using the proposed AD or SD, depending on the number of layers $L$. Specifically, we want to illustrate the results presented in Theorem 2.5 and in Theorem 3.8. Although these results have only been shown for $L=1$ or $L\to\infty$ (i.e., the proximity operators are computed accurately), we will also empirically investigate the behavior of the AD and SD when $1<L<+\infty$. To this aim, we consider a toy example consisting of an equation system of the form of $y=A\overline{x}$, where $\overline{x}\in\mathbb{R}^{N}$ is a vector of dimension $N=50$ generated randomly following a uniform distribution, $A\colon\mathbb{R}^{N}\to\mathbb{R}^{M}$ is a Gaussian measurement operator with $M=20$, where coefficients are generated randomly following a Gaussian distribution with mean 0 and standard deviation $1$. Both the analysis and synthesis dictionaries are generated as Gaussian dictionaries with mean 0 and standard deviation $1$, of size $\Gamma\in\mathbb{R}^{100\times 50}$ and $D\in\mathbb{R}^{50\times 100}$, respectively.
@@ 5.1.1 Analysis denoiser
Figure 1: Convergence profile of algorithm (2.7) with $L\in\{1,20,50,100\}$. Solutions $(x^{*},u^{*})$ are pre-computed considering algorithm (2.7) with $L\to\infty$.
@@
In Figure 1 we show the behavior of $(\|x_{k}-x^{*}\|)_{k\in\mathbb{N}}$ and $(\|u_{k}-u^{*}\|)_{k\in\mathbb{N}}$ with $(x_{k},u_{k})_{k\in\mathbb{N}}$ generated by algorithm (2.7) with $L\in\{1,20,50,100\}$, and $(x^{*},u^{*})$ generated by algorithm (2.7) with $L\to\infty$ (obtained after $10^{4}$ iterations). Note that $x^{*}$ is solution to problem (2.6). We see that the trajectory per iteration is the same independently of the value of $L$. Hence, according to Theorem 2.5 we can deduce that $(x^{*},\tau^{-1}u^{*})$ is solution to the saddle-point problem (2.9). We can further observe that the convergence is faster in time for smaller values of $L$.
@@ 5.1.2 Synthesis denoiser
We showed in Theorem 3.8 that taking $L=1$ in algorithm (3.20) enables solving (3.6) as well as the direct sparse coding problem (3.7). Since algorithm (3.20) for $L\to\infty$ solves (3.6) (Theorem 3.7), we deduce that both algorithm (3.20) for $L\to\infty$ and algorithm (3.21) (i.e., $L=1$) solve the same problem (3.6). In this section we aim to illustrate this result by comparing the output of algorithm (3.21) with the output of algorithm (3.20) for $L=10^{4}$. Furthermore, although the results from Proposition 3.1 only hold for $L=1$, we show empirically that they remain true for $1<L<+\infty$ for our problem by considering $L\in\{20,50,100\}$.
In Figure 2 we show the behavior of $(\|x_{k}-x^{*}\|)_{k\in\mathbb{N}}$ and $(\|z_{k}-z^{*}\|)_{k\in\mathbb{N}}$ with $(x_{k},z_{k})_{k\in\mathbb{N}}$ generated by algorithm (3.20) with $L\in\{1,20,50,100\}$, and $(x^{*},z^{*})$ generated by algorithm (3.20) with $L\to\infty$ (obtained after $10^{4}$ iterations). Note that $x^{*}$ is solution to problem (3.6). We see that the trajectory per iteration is fairly the same when $L\in\{20,50,1000\}$, while $L=1$ is taking a bit more iterations to reach convergence. However, for the convergence speed, all strategies seems to behave similarly, with $L\in\{1,20\}$ being very slightly faster.
Figure 2: Convergence profile of algorithm (3.20) with $L\in\{1,20,50,100\}$. Solutions $(x^{*},u^{*})$ are pre-computed considering algorithm (3.20) with $L\to\infty$.
@@ 5.2 Application to image deblurring within a DDL framework
We now consider an image restoration problem to illustrate results presented in Sections 2 and 3. Although the results presented in this work are generic and can be applied for any linear operators $\Gamma$ and $D$, they are of particular interest in the DDL framework, where these operators are learned. We will explore this case in this section, and apply the resulting methods to an image deblurring problem. For the sake of completeness we will compare the resulting algorithms to a FB-PnP algorithm where the denoiser is a DRUnet.
@@ 5.2.1 DDL context and training
$v=\overline{x}+{\varepsilon}w,$ (5.1)
where $w\in\mathbb{R}^{N}$ is a realization of an i.i.d. standard normal random variable, and ${\varepsilon}>0$ is the noise level. Both analysis and synthesis formulations to find the estimate of $\overline{x}$ from $v$ in (5.1) (see (1.6) and (3.1), respectively) have been extensively studied in the literature (see e.g. [46, 60] for the analysis, and [44, 2, 39] for the synthesis formulation). In the context of DDL, these problems can be reformulated as a joint estimation of the image and the dictionary, or as bi-level optimization problems [39]. Then, the dictionaries are defined as solutions to
@@
If in (5.2) (resp. (5.3)), the loss function $\ell^{\mathcal{A}}(\Gamma)$ is the same as the one minimized in (1.6) (resp. if $\ell^{\mathcal{S}}(D)$ is the same as the one minimized in (1.7)), then the bi-level problem is equivalent to the DDL joint estimation problem. This choice is often used in an unsupervised setting. In a supervised setting, where a ground truth dataset can be used, a standard choice for $\ell^{\mathcal{A}}$ and $\ell^{\mathcal{S}}$ is the mean square error (MSE) loss, minimizing the $\ell^{2}$ norm of the difference between the ground truth and the estimated signal $x^{\dagger}_{\Gamma}$ or $x^{\dagger}_{D}$. In the remainder, we will use this approach. Specifically, we aim to learn $\Gamma$ and $D$ such that, for some initializations $u_{0}\in\mathbb{R}^{S}$ and $z_{0}\in\mathbb{R}^{S}$, we have
$\begin{cases}x^{\dagger}_{\mathcal{A}}=G^{\mathcal{A}}_{L,\lambda,v}({u_{0}})\approx\overline{x},\\ x^{\dagger}_{\mathcal{S}}=G^{\mathcal{S}}_{L,\lambda,v}({z_{0}})\approx\overline{x}\end{cases}$ (5.4)
@@
where $G^{\mathcal{A}}$ and $G^{\mathcal{S}}$ are defined in Model 2.2 and Model 3.5, respectively, for a fixed number of iterations $L\geq 1$, a regularization parameter $\lambda>0$, and input noisy image $v$ obtained as per (5.1). We assume that dictionaries $\Gamma$ and $D$ are parametrized by learnable parameters $\theta_{\Gamma}$ and $\theta_{D}$, respectively (i.e., convolution kernel). Then, given a dataset of pairs of clean/noisy images ${(\overline{x}_{i},v_{i})}_{i\in\mathbb{S}_{\mathcal{T}}}$ obtained following (5.1), the unrolled denoisers in (5.4) are trained by solving, respectively
@@
$\underset{\theta_{\Gamma}}{\operatorname{minimize}}\;\frac{1}{\#(\mathbb{S}_{\mathcal{T}})}\!\sum_{i\in\mathbb{S}_{\mathcal{T}}}\!\|G^{\mathcal{A}}_{L_{\text{train}},\lambda,v_{i}}(0_{S})-\overline{x}_{i}\|,\,\text{ and }\,\underset{\theta_{D}}{\operatorname{minimize}}\;\frac{1}{\#(\mathbb{S}_{\mathcal{T}})}\!\sum_{i\in\mathbb{S}_{\mathcal{T}}}\!\|G^{\mathcal{S}}_{L_{\text{train}},\lambda,v_{i}}(0_{S})-\overline{x}_{i}\|,$
[…]
@@
Denoising performances – In Table 3 we give the inference time for denoising an image, when applying the AD and SD with either $L_{\text{den}}=1$ or $L_{\text{den}}=20$ iterations. For comparison, we also provide the inference time when using a DRUnet, where the weights have been taken from the DeepInv PyTorch library. We further provide examples of output images for the different denoisers in Figure 3, evaluated in different configurations. It can be observed that the cases where SD is evaluated with $L_{\text{den}}=1$ iteration (SD train $L_{\text{train}}=\{1,20\}$ and $L_{\text{den}}=1$) lead to very poor results. When SD is trained with $L_{\text{train}}=1$ and evaluated with $L_{\text{den}}=\{20,1000\}$ iterations, results start to improve, although remaining quite poor. All other cases look very similar, with DRUnet giving best denoising results.
@@
Table 3: Average time and standard deviation when applying different denoisers. Second row shows time to evaluate the denoisers with either $L_{\text{den}}=1$ or $L_{\text{den}}=20$, computed over $150$ images from BSDS500 cropped to $150\times 150$, with noise level ${\varepsilon}=0.05$. Third row shows time to run $1,000$ iterations of the PnP with either $L_{\text{PnP}}=1$ or $L_{\text{PnP}}=20$, computed over 44 examples.
@@
Average time per evaluation $0.19(\pm 0.08)$ $3.94(\pm 1.46)$ $0.27(\pm 0.16)$ $3.12(\pm 0.96)$ $5.32(\pm 0.76)$
[…]
Average time for PnP run $1.32(\pm 0.03)$ $31.07(\pm 0.06)$ $1.18(\pm 0.02)$ $18.57(\pm 0.06)$ $49.24(\pm 0.19)$
[…]
@@
$\overline{x}$ $L_{\text{den}}=1$ $L_{\text{den}}=20$ $L_{\text{den}}=1$ $L_{\text{den}}=20$
@@
$v$ ($\varepsilon=0.05$) $L_{\text{den}}=1$ $L_{\text{den}}=20$ $L_{\text{den}}=1$ $L_{\text{den}}=20$
@@
Denoising for $L_{\text{den}}=1,000$
DRUnet SD ($L_{\text{train}}=1$) SD ($L_{\text{train}}=20$) AD ($L_{\text{train}}=1$) AD ($L_{\text{train}}=20$)
35.835.8dB 26.726.7dB 26.626.6dB 29.029.0dB 31.431.4dB
Figure 3: Example of outputs obtained with the unfolded SD and AD for different configurations. Left column shows original image $\overline{x}$, noisy image $v$ and denoised image obtained with DRUnet for comparison. Other columns are as follows. First (resp. second) row: Outputs of the SD (resp. AD) for different settings $(L_{\text{train}},L_{\text{den}})\in\{1,20\}^{2}$. Last row: Denoised images obtained for the SD and the AD for the two training settings $L_{\text{train}}\in\{1,20\}$ and with $L_{\text{den}}=1,000$.
@@ 5.2.2 Simulation results
We consider an image deblurring problem of the form of (1.1), where $A\colon\mathbb{R}^{N}\to\mathbb{R}^{M}$ models a convolution with kernel given in Figure 4 and $\varepsilon=0.05$. Both the AD and SD are trained following the DDL procedure described in Section 5.2.1. The resulting AD and SD are then plugged in the FB-PnP iterations and applied to the image deblurring problem. We will further compare the proposed approaches with an FB-PnP algorithm with a DRUnet denoiser [61], considered as state-of-the-art for PnP (see Section B).
@@
Figure 5: Convergence profiles with respect to iterations for the different FB-PnP algorithms. AD and SD are trained with $L_{\text{train}}\in\{1,20\}$, and implemented in the FB-PnP with either $L_{\text{PnP}}\in\{1,20\}$.
@@
Figure 6: PSNR profiles with respect to iterations for the different FB-PnP algorithms. AD and SD are trained with $L_{\text{train}}\in\{1,20\}$, and implemented in the FB-PnP with either $L_{\text{PnP}}\in\{1,20\}$.
@@
In Figure 5, we compare convergence behaviors of the FB-PnP iterations $(\|x_{k+1}-x_{k}\|)_{k\in\mathbb{N}}$ for different denoisers, for different regularization parameters $\lambda\in\{10^{-3},3^{-3},10^{-2},3^{-2},10^{-1}\}$. Each plot shows convergence curves when running the algorithm on $4$ different images. The first (resp. second) row shows results obtained with SD (resp. AD), trained on different settings $L_{\text{train}}\in\{1,20\}$ and evaluated on different settings $L_{\text{PnP}}\in\{1,20\}$. The third row in Figure 5 shows FB-PnP behaviour with DRUnet. FB-PnP algorithms with SD and AD exhibit monotonic convergence trends independently of the choice of the regularization parameter $\lambda$. This is not the case for DRUnet that seems to converge only for $\lambda=10^{-1}$, while other choices are not stable [^1] For DRUnet $\lambda$ corresponds to the noise level given as an input to the network (see Section B for details). Additional simulations have been performed for DRUnet increasing $\lambda$, however these choices were observed to lead to lower PSNR values. . We further provide curves of PSNR values with respect to iterations in Figure 6, for the above mentioned FB-PnP iterations. It can be observed that for AD the PSNR profiles are very similar when fixing $\lambda$, and looking at the different strategies $(L_{\text{train}},L_{\text{PnP}})\in\{1,20\}^{2}$. This suggests that the AD framework is stable, independently on the number of sub-iterations used for training or for the reconstruction process. This is not the case for SD. We can observe that the strategies $(L_{\text{train}},L_{\text{PnP}})=(1,20)$ and $(L_{\text{train}},L_{\text{PnP}})=(20,20)$ have similar trends, and strategies $(L_{\text{train}},L_{\text{PnP}})=(1,1)$ and $(L_{\text{train}},L_{\text{PnP}})=(20,1)$ have similar trends too. The last two strategies seems also to convergence slower. This was to be expected given theirs denoising performances shown in Figure 3 for $(L_{\text{train}},L_{\text{den}})=(1,1)$ and $(L_{\text{train}},L_{\text{den}})=(20,1)$. DRUnet exhibits very unstable behaviour for $\lambda\in\{10^{-3},3\times 10^{-3},10^{-2}\}$, as expected given Figure 5. More stable results are for $\lambda=10^{-1}$. The choice $\lambda=3\times 10^{-2}$ seems to give slightly higher PSNR values, although the associated convergence profiles in Figure 5 are not very stable (no monotony as per convergence guarantees).
@@
Figure 7: Examples of FB-PnP outputs obtained with the different denoisers tested, for two image examples. Regularization parameters are chosen according to convergence profiles given in Figure 5 and Figure 6: $\lambda_{\textsf{SD}}=3\times 10^{-2}$ for $(L_{\text{train}},L_{\text{PnP}})\in\{(1,1),(20,1),(20,20\}$, $\lambda_{\textsf{SD}}=10^{-2}$ for $(L_{\text{train}},L_{\text{PnP}})=(1,20)=10^{-10}$, $\lambda_{\textsf{AD}}=3\times 10^{-3}$ for all $(L_{\text{train}},L_{\text{PnP}})$ and $\lambda_{\textsf{DRUnet}}=10^{-1}$.